# Isomorphism classes of generating sets

### v1

@@ Abstract.
We prove that for any two regular cardinals $\omega<\lambda_{0}<\lambda_{1}$ there is a ccc forcing extension where there is an ultrafilter $U$ on $\omega$ with a base $\mathcal{B}$ such that $(\mathcal{B},\supseteq^{*})\cong\lambda_{0}\times\lambda_{1}$. We use similar ideas to construct an ultrafilter with a base $\mathcal{B}$ as above which is order isomorphic to any given two-dimensional, well-founded, countably directed order with no maximal element. Similarly, relative to a supercompact cardinal, it is consistent that $\kappa$ is supercompact, and for any regular cardinals $\kappa<\lambda_{0}<\lambda_{1}<...<\lambda_{n}$, there is a ${<}\kappa$-directed closed $\kappa^{+}$-cc forcing extension where there is a normal ultrafilter $U$ on $\kappa$ with a base $\mathcal{B}$ such that $(\mathcal{B},\supseteq^{*})\cong\lambda_{0}\times...\times\lambda_{n}$. We apply our constructions to obtain ultrafilters with controlled Tukey-type, in particular, an ultrafilter with non-convex Tukey and depth spectra is presented, answering questions from [4]. Our construction also provides new models where $\mathfrak{u}_{\kappa}<2^{\kappa}$.
@@
The results of this paper were conceived at the American Institute of Mathematics in Caltech as part of a SQuaRE program.
@@ 1. Introduction
Ultrafilters with strong combinatorial properties have been proven useful for the construction of mathematical objects with extreme behavior such as ultraproducts, topological spaces (and special points in topological spaces), and certain groups, but could also be used to prove some combinatorial properties [32, 28, 34]. Many of these combinatorial properties are obtained by constructing a special generating set for the ultrafilter. This paper aims to study the possible structures of such generating sets. More precisely, we are interested in the isomorphism class of the structure $(\mathcal{B},\supseteq^{*})$, where $\supseteq^{*}$ is the reversed inclusion modulo the bounded ideal (or finite in case of $\omega$) and $\mathcal{B}$ is a $\subseteq^{*}$-generating set of an ultrafilter $U$, namely, $\mathcal{B}\subseteq U$ and for every $A\in U$, there is $B\in\mathcal{B}$ such that $B\subseteq^{*}A$.
@@
Assume GCH and let $\lambda_{0},\lambda_{1}$ be cardinals of uncountable cofinality. Then there is a ccc forcing extension in which there is an ultrafilter $U$ on $\omega$, with a $\subseteq^{*}$-generating set $\mathcal{B}\subseteq U$ such that $(\mathcal{B},\supseteq^{*})\simeq\lambda_{0}\times\lambda_{1}$.
The proof uses non-linear iterations of Mathias forcing. Such iterations were considered before by Hechler [29] who proved a similar result to obtain dominating families in the structure $(\omega^{\omega},\leq^{*})$. Several results in the spirit of Hechler can be found in the literature [23, 17, 1]. Blass and Shelah [13] considered a non-linear iteration involving Mathias forcing, however, the Mathias component only really appears in one dimension of this construction. The major difficulty, however, is to obtain at the end of the process (i.e. iteration) a uniform ultrafilter.
@@
[…]
@@
Suppose that $\mathbb{D}$ is a well-founded, two-dimensional, countably directed lattice with no maximal element, then there is a ccc extension in which there is a simple $P_{\mathbb{D}}$-point ultrafilter on $\omega$.
Generalizing Kunen’s result to measurable cardinals requires considerable effort and does not run smoothly using the standard indestructibility methods. In an unpublished work, Carlson showed that from a supercompact cardinal this is possible. Recently, this type of iteration appeared in several constructions [27, 18, 16, 22] guided by (for example) a diamond sequence so that a supercompact embedding lifts in a way that the normal ultrafilter derived from the lifted embedding has a $\subseteq^{*}$-decreasing generating set of the desired length. This can be used to obtain a model with $2^{\kappa}>\kappa^{+}$, with a $\kappa$-complete ultrafilter over $\kappa$ generated by fewer than $2^{\kappa}$-many sets. It remains open whether this can be achieved from weaker assumptions.
Our results also apply to this type of lifting argument. In particular, we obtain the following:
Theorem 1.4.
Relative to the existence of a supercompact cardinal, there is a model $V$ with a supercompact cardinal $\kappa$ such that for every regular cardinals $\kappa<\lambda_{0}<\lambda_{1}$, there is a ${<}\kappa$-directed closed $\kappa^{+}$-cc forcing extension in which there is a $\kappa$-complete simple $P_{\lambda_{0}\times\lambda_{1}}$-point ultrafilter on $\kappa$.
Again, we can use general order theoretic reductions to obtain further dimension-two orders as simple $P$-point ultrafilters. However, while the construction on $\omega$ does not work for three-dimensional orders (such as $\lambda_{0}\times\lambda_{1}\times\lambda_{2}$), the diamond mechanism is more flexible in that sense and in fact we obtain the following generalization:
Theorem 1.5.
Relative to the existence of a supercompact cardinal, there is a model $V$ with a supercompact cardinal $\kappa$ such that for all regular cardinals $\kappa<\lambda_{0}<\lambda_{1}<...<\lambda_{n}$, there is a ${<}\kappa$-directed closed $\kappa^{+}$-cc forcing extension in which there is a $\kappa$-complete simple $P_{\lambda_{0}\times\lambda_{1}\times...\times\lambda_{n}}$-point ultrafilter over $\kappa$.
These results can also be used to construct complex generating sets for the club filter on a measurable cardinal:
@@
Relative to the existence of a supercompact cardinal, there is a model $V$ with a supercompact cardinal $\kappa$ such that for every regular cardinals $\kappa<\lambda_{0}<\lambda_{1}<...<\lambda_{n}$, there is a ${<}\kappa$-directed closed $\kappa^{+}$-cc forcing extension in which $\kappa$ is measurable and the club filter is a simple $P_{\lambda_{0}\times...\times\lambda_{n}}$-point.
@@
Finally, we apply these constructions to control the Tukey-type of certain ultrafilters. Recall that the Tukey order on directed sets is defined by $(\mathbb{P},\leq_{P})\leq_{T}(\mathbb{Q},\leq_{Q})$ if there is a Tukey reduction $f:\mathbb{P}\to\mathbb{Q}$, that is, for every unbounded $\mathcal{B}\subseteq\mathbb{P}$, $f[\mathcal{B}]$ is unbounded in $\mathbb{Q}$. We say that $\mathbb{P}\equiv_{T}\mathbb{Q}$ if $\mathbb{P}\leq_{T}\mathbb{Q}$ and $\mathbb{Q}\leq_{T}\mathbb{P}$. In this paper, we will only consider ultrafilters $U$ ordered by reversed inclusion $(U,\supseteq)$ and almost inclusion $(U,\supseteq^{*})$. The study of the Tukey order find its origins in the concept of Moore-Smith convergence of nets and has been studied extensively on general ordered sets [46, 30, 44]. The Tukey-type $(U,\supseteq)$ also has been subject to a considerable amount of work, especially when $U$ is an ultrafilter on $\omega$ [30, 19, 20, 12, 38, 39, 36, 11, 3] and lately this has been also looked at in the context of measurable cardinals [5, 9] due to its close relation to the Galvin property and applicability to Prikry-type forcing theory [8, 26, 6]. The Tukey-type of $(U,\supseteq^{*})$ as also been of interest [36, 5] as the properties of a $\subseteq^{*}$-generating set for an ultrafilter are related to this Tukey-type. In fact, most of the combinatorial characteristics of ultrafilters (see for example [15]) are formulated in terms of $\subseteq^{*}$ rather than $\subseteq$. The connection to the Tukey order is noticeable once the Tukey order is expressed in-terms of cofinal maps. Given a partially ordered set $(D,\leq_{D})$, a set $B\subseteq D$ is called cofinal in $(D,\leq_{D})$ if for every $d\in D$ there is $b\in B$ such that $d\leq_{D}b$. Schmidt Duality [40] in the special case of ultrafilters (or in the situation where there are least upper bounds): For any two ultrafilters $U,W$, $U\leq_{T}W$ if and only if there is a function $f:W\to U$ which is monotone (i.e. $A\subseteq B\Rightarrow f(A)\subseteq f(B)$) and has a cofinal image in $U$.
@@
It is not hard to check that whenever $U$ is a simple $P_{\mathbb{D}}$-point, then $(U,\supseteq^{*})\equiv_{T}\mathbb{D}$. So, the analysis in this paper is finer than the Tukey-type analysis of ultrafilters, but can be used to control it. For example, we use our construction to find an ultrafilter $U$ which is Tukey-equivalent to $\lambda_{0}\times\lambda_{1}$. From this we can analyze the Tukey spectrum and the depth spectrum (see Definitions 5.1,  5.6) and show the consistency of an ultrafilter with a non-convex spectrum (both depth and Tukey). This was asked by the first author in [4].
@@
[…]
@@
As we mentioned before, at a measurable cardinal, the only known method to produce models where $\mathfrak{u}_{\kappa}<2^{\kappa}$ is via the linear Mathias iteration. The computation of the values of these cardinal characteristics was done in [16] where it was shown that after a linear iteration of Mathias forcing, $\mathfrak{u}_{\kappa}=\mathfrak{u}^{com}_{\kappa}=\mathfrak{b}_{\kappa}=\mathfrak{d}_{\kappa}$. Lately, the first and the third author [10] generalized this and showed that merely the presence of a simple $P_{\lambda}$-point (which is the kind of ultrafilters produced by the Mathias iteration) causes $\mathfrak{u}_{\kappa}=\mathfrak{u}^{com}_{\kappa}=\mathfrak{b}_{\kappa}=\mathfrak{d}_{\kappa}$. So the question of separating the cardinal characteristics in a model where $\mathfrak{u}_{\kappa}<2^{\kappa}$ remains open. Using our methods, we can prove the following:
@@ 2. Some preliminaries
Given a partial order $(D,\leq_{D})$, we define $d_{1}<_{D}d_{2}$ if $d_{1}\leq_{D}d_{2}$ and $d_{1}\neq d_{2}$. For each $d\in D$, let $D_{<d}=\{e\in D:e<_{D}d\}$; $D_{\leq d}$ is defined similarly.
@@
[…]
@@
Since $X_{\text{even}}\cup X_{\text{odd}}=A^{*}\in U$, either $X_{\text{even}}\in U$ or $X_{\text{odd}}\in U$. Suppose for example that $X_{\text{even}}\in U$ (the argument in the case that $X_{\text{odd}}\in U$ is identical). Then there is $i_{1}<\kappa$ such that $X_{\text{even}}\in F_{i_{1}}$. By $(2)$ of the claim, and $\kappa$-completeness of $F_{i_{1}}$, $\bigcap_{\alpha<i_{1}}A_{\alpha}\in F_{i_{1}}$. Therefore $X_{\text{even}}\cap\bigcap_{\alpha<i_{1}}A_{\alpha}\in F_{i_{1}}$. Since the $F_{r}$’s are increasing, we may assume that $i_{1}=2\gamma+1$ for some $\gamma<\kappa$. Finally, we claim that $X_{\text{even}}\cap\bigcap_{\alpha<2\gamma+1}A_{\alpha}\subseteq A_{2\gamma+1}$, which produces the desired contradiction. Indeed if $\nu\in X_{\text{even}}\cap\bigcap_{\alpha<2\gamma+1}A_{\alpha}$ then by definition of $X_{\text{even}}$ there is $\alpha^{*}<\kappa$ such that $\nu\in C_{2\alpha^{*}}$. Since $\nu\in\bigcap_{\alpha<2\gamma+1}A_{\alpha}$, it follows that $12\alpha^{*}>2\gamma+$ 1. This means that $\nu\in A_{2\gamma+1}.$
@@
The ultrafilters we will be constructing are all $p$-points. This poses more restrictions on the possible $\mathbb{D}$ which we will be able to realize. Recall that an ultrafilter $U$ is a $P_{\lambda}$-point, if $(U,\supseteq^{*})$ is a $\lambda$-directed ordered set.
@@
We now turn to the order theoretic part of the preliminaries. The following definition will be essential in our construction of a simple $P_{\mathbb{D}}$-point with a prescribed $\mathbb{D}$.
Definition 2.6.
[…]
Given a lattice $\mathbb{D}$, a sublattice $\mathbb{A}$ of $\mathbb{D}$ is a suborder of $\mathbb{D}$ which is closed under joins and meets.
Example 2.7.
Any product of ordinals $\lambda_{0}\times\lambda_{1}$ is a lattice of degree $2$: Given pairs $(\alpha_{1},\beta_{1}),(\alpha_{2},\beta_{2}),(\alpha_{3},\beta_{3})<(\alpha,\beta)$,
$\alpha_{\max}=\max_{1\leq i\leq 3}\alpha_{i},\ \alpha_{\min}=\min_{1\leq i\leq 3}\alpha_{i},\ \beta_{\max}=\max_{1\leq i\leq 3}\beta_{i},\ \beta_{\min}=\min_{1\leq i\leq 3}\beta_{i}.$
Then
$(\alpha_{1},\beta_{1})\vee(\alpha_{2},\beta_{2})\vee(\alpha_{3},\beta_{3})=(\alpha_{\max},\beta_{\max})$
$(\alpha_{1},\beta_{1})\wedge(\alpha_{2},\beta_{2})\wedge(\alpha_{3},\beta_{3})=(\alpha_{\min},\beta_{\min}).$
If the pairs are pairwise incomparable, then for each $i=1,2,3$, either $\alpha_{i}<\alpha_{\max}$ or $\beta_{i}<\beta_{\max}$. By the pigeonhole principle, there are $i\neq j$ such that either $\max\{\alpha_{i},\alpha_{j}\}<\alpha_{\max}$ or $\max\{\beta_{i},\beta_{j}\}<\beta_{\max}$. Without loss of generality suppose that $\max\{\alpha_{i},\alpha_{j}\}<\alpha_{\max}$. This means that
$(\max(\alpha_{i},\alpha_{j}),\max(\beta_{i},\beta_{j}))<(\alpha_{\max},\beta_{\max})\leq(\alpha,\beta).$
Hence $\lambda_{0}\times\lambda_{1}$ has degree $2$.
Proposition 2.8.
  1. (1)
    Suppose that $\mathbb{D}$ is a degree $2$ lattice and that $\mathbb{A}\subseteq\mathbb{D}$ is a sub-lattice. Then $\mathbb{A}$ is a degree $2$ lattice.
  2. (2)
    If $\mathbb{D}^{\prime}$ is a directed set which embeds into a degree $2$ lattice $\mathbb{D}$, then it also embeds cofinally into a degree $2$ sub-lattice.

### v2

@@ Abstract.
We introduce a new class of ultrafilters which generalizes the well-known class of simple $P$-point ultrafilters. We prove that for any well-founded $\sigma$-directed partial order $\mathbb{D}$ there is a mild forcing extension where there is an ultrafilter $U$ on $\omega$ with a base $\mathcal{B}$ such that $(\mathcal{B},\supseteq^{*})\cong\mathbb{D}$. On a measurable cardinal we prove a similar result: relative to a supercompact cardinal, it is consistent that $\kappa$ is supercompact, and for a $\kappa^{+}$-directed well-founded poset $\mathbb{D}$, there is a $\kappa-$directed closed $\kappa^{+}$-cc forcing extension where there is a normal ultrafilter $U$ on $\kappa$ with a base $\mathcal{B}$ such that $(\mathcal{B},\supseteq^{*})\cong\mathbb{D}$. These are optimal results in the class of $P$-points and realize every potential structure of a $P$-point. We apply our constructions to obtain ultrafilters with controlled Tukey-type, in particular, an ultrafilter with non-convex Tukey and depth spectra is presented, answering questions from [10]. Our construction also provides new models where $\mathfrak{u}_{\kappa}<2^{\kappa}$, answering questions from [7].
Key words and phrases:
Simple p-point, ultrafilter number, Tukey order, supercompact cardinals, non-linear iterated forcing
2010 Mathematics Subject Classification:
03E04, 06A07, 54D80, 03E55
@@
The forth author was partially supported by the Israel Science Fund, ISF 1967/21.
@@
The results of this paper were conceived at the American Institute of Mathematics in Pasadena as part of a SQuaRE program. We thank the Institute for their support and for providing an excellent working environment.
@@ 1. Introduction
Ultrafilters with strong combinatorial properties have proven useful in constructing mathematical objects exhibiting extreme behavior, such as ultraproducts, topological spaces (including special points within them), and certain groups. They can also be employed to establish purely combinatorial results [32, 28, 33]. Many of these results are obtained by constructing a special generating set for the ultrafilter. This paper aims to study the possible structures of such generating sets. More precisely, we are interested in the isomorphism class of the structure $(\mathcal{B},\supseteq^{*})$, where $\supseteq^{*}$ is the reversed inclusion modulo the bounded ideal (or finite in case of $\omega$) and $\mathcal{B}$ is a $\subseteq^{*}$-generating set of a filter $F$, namely, $\mathcal{B}\subseteq F$ and for every $A\in F$, there is $B\in\mathcal{B}$ such that $B\subseteq^{*}A$.
@@
[…]
The proof uses non-linear iterations of Mathias forcing. Such iterations were considered before by Hechler [29] who proved a similar result to obtain dominating families in the structure $(\omega^{\omega},\leq^{*})$. Several results in the spirit of Hechler can be found in the literature [22, 17, 1]. Blass and Shelah [13] considered a non-linear iteration involving Mathias forcing, however, the Mathias component only really appears in one dimension of this construction. The major difficulty, however, is to obtain at the end of the process (i.e. iteration) a uniform ultrafilter. Milovich [37, Thm. 5.10], produced via traditional linear iteration an ultrafilter $U$ on $\omega$ which is generated by a set $\mathcal{B}$ isomorphic to a cofinal subset of $\lambda_{0}\times\lambda_{1}$.
@@
Given an ordered set $\mathbb{D}=(D,\leq_{D})$, we say that a filter $F$ is a simple $P_{\mathbb{D}}$-point if there is a $\subseteq^{*}$-generating set $\mathcal{B}\subseteq F$ such that $(\mathcal{B},\supseteq^{*})\simeq\mathbb{D}$.
Hence Theorem 1.1 says that in a forcing extension, there is a uniform simple $\mathbb{P}_{\lambda_{0}\times\lambda_{1}}$-point ultrafilter. Our result is much more general and applies to every well-founded $\sigma$-directed set.
Theorem 1.3.
@@
Starting from any model, the $\lambda^{+}$-closed forcing in the previous theorem is simply the standard forcing to add a $\diamondsuit_{\lambda^{+}}(\lambda^{+}\cap\text{cof}(\lambda))$ sequence with initial segments. If this diamond already holds in the ground model, we can directly force with the ccc forcing. We also note that our method works equally well if $\diamondsuit_{\mu}(\mu\cap\text{cof}(\lambda))$ holds for any $\mu\geq\lambda^{+}$ regular. Hence we have the following corollary:
Corollary 1.4.
Let $\mathbb{D}$ be a well-founded, $\sigma$-directed set, which has a possible order type $\lambda$. Assume that $\diamondsuit_{\mu}(\mu\cap\text{cof}(\lambda))$ holds for some regular cardinal $\mu>\lambda$. There is a ccc forcing extension in which $2^{\omega}=\mu$ and there is a simple $P_{\mathbb{D}}$-point ultrafilter on $\omega$.
It is essential to note that for any given $\mathbb{D}$, a simple $\mathbb{P}_{\mathbb{D}}$-point can easily be forced if we do not require some nice properties of the forcing which is used to create it, so as to preserve the properties of $\mathbb{D}$. For example, we can always collapse $|\mathbb{D}|$ to $\omega_{1}$ and run Kunen’s argument. Hence the strength of Theorem 1.3 and Corollary 1.4 lies within the type of forcings we use. In Section 4.2 we prove these theorems, and at the beginning of Section 5 we establish that the chain condition and closure of the forcing ensures the preservation of the structure of $\mathbb{D}$.
Generalizing Kunen’s result to measurable cardinals requires considerable effort and does not run smoothly using the standard indestructibility methods. In an unpublished work, Woodin and Carlson showed that from a supercompact cardinal, this is possible. Recently, this type of iteration appeared in several constructions [25, 18, 16, 21] guided by (for example) a diamond sequence so that a supercompact embedding lifts in a way that the normal ultrafilter derived from the lifted embedding has a $\subseteq^{*}$-decreasing generating set of the desired length. This can be used to obtain a model with $2^{\kappa}>\kappa^{+}$, with a $\kappa$-complete ultrafilter over $\kappa$ generated by fewer than $2^{\kappa}$-many sets. It remains open whether this can be achieved from weaker large cardinal assumptions. It is this type of argument which we use to prove Theorem 1.3.
Our method is flexible enough to incorporate lifting arguments and produce similar results on measurable cardinals:
@@
Let $\kappa$ be a supercompact cardinal and $\mathbb{D}$ be a well-founded, $\kappa^{+}$-directed set with a possible order type $\lambda$. Then for every regular $\mu>\lambda$, there is an extension by a ($\kappa$-$c$) $\ast$ ($\mu$-directed-closed) $\ast$ ($\kappa$-directed-closed $\kappa^{+}$-cc) forcing in which there is a normal ultrafilter which is a simple $P_{\mathbb{D}}$-point ultrafilter over $\kappa$.
The first forcing here is the the usual Laver preparation [35] which makes the supercompactness of $\kappa$-indestructible under $\kappa$-directed closed forcings, and the second forcing adds a $\diamondsuit_{\mu}(\mu\cap\text{cof}(\lambda))$-sequence. Since the Laver preparation forcing is $\kappa$-cc with cardinality $\kappa$, it will preserve such a diamond sequence if it already existed.
@@
Let $\kappa$ be a supercompact cardinal, let $\mathbb{D}$ be a well-founded, $\kappa^{+}$-directed set with a possible order type $\lambda$, and let $\mu>\lambda$ be a regular cardinal such that $\diamondsuit_{\mu}(\mu\cap\text{cof}(\lambda))$ holds. Then there is an extension by a ($\kappa$-$c$) $\ast$ ($\kappa$-directed-closed $\kappa^{+}$-cc) forcing in which there is a normal ultrafilter which is a simple $P_{\mathbb{D}}$-point ultrafilter over $\kappa$.
These results can also be used to construct complex generating sets for the club filter on a measurable cardinal as well:
Corollary 1.7.
Let $\kappa$ be a supercompact cardinal and $\mathbb{D}$ be a well-founded, $\kappa^{+}$-directed set with a possible order type $\lambda$. Then for every regular $\mu>\lambda$, there is an extension by a ($\kappa$-$c$) $\ast$ ($\mu$-directed-closed) $\ast$ ($\kappa$-directed-closed $\kappa^{+}$-cc) forcing in which $\kappa$ is measurable and the club filter on $\kappa$ is a simple $P_{\mathbb{D}}$-point.
Theorem 1.5 and Corollaries 1.6,1.7 will be proven in Section 4.1.
All the ultrafilters which are produced in this paper are $P$-points. This is an integral feature of the Mathias iteration. Our results are optimal in the class of $P$-points since if a $P$-point over $\kappa\geq\omega$ is a simple $\mathbb{P}_{\mathbb{D}}$-point, then $\mathbb{D}$ has to be $\kappa^{+}$-directed. It is also evident that every order has a well-founded cofinal subset. So our results say that every such order can also be realized.
@@
Finally, in Seciton 5, we provide several applications of our construction. The first is the ability to control the Tukey-type of certain ultrafilters. Recall that the Tukey order on directed sets is defined by $(\mathbb{P},\leq_{P})\leq_{T}(\mathbb{Q},\leq_{Q})$ if there is a Tukey reduction $f:\mathbb{P}\to\mathbb{Q}$, that is, for every unbounded $\mathcal{B}\subseteq\mathbb{P}$, $f[\mathcal{B}]$ is unbounded in $\mathbb{Q}$. We say that $\mathbb{P}\equiv_{T}\mathbb{Q}$ if $\mathbb{P}\leq_{T}\mathbb{Q}$ and $\mathbb{Q}\leq_{T}\mathbb{P}$. In this paper, we will only consider ultrafilters $U$ ordered by reversed inclusion $(U,\supseteq)$ and almost inclusion $(U,\supseteq^{*})$. The study of the Tukey order finds its origins in the concept of Moore-Smith convergence of nets and has been studied extensively on general ordered sets [47, 30, 46]. The Tukey-type $(U,\supseteq)$ also has been subject to a considerable amount of work, especially when $U$ is an ultrafilter on $\omega$ [30, 20, 19, 12, 41, 42, 38, 8, 9] and lately this has been also looked at in the context of measurable cardinals [2, 6] due to its close relation to the Galvin property and applicability to Prikry-type forcing theory [5, 26, 3]. The Tukey-type of $(U,\supseteq^{*})$ as also been of interest [38, 2] as the properties of a $\subseteq^{*}$-generating set for an ultrafilter are related to this Tukey-type. In fact, most of the combinatorial characteristics of ultrafilters (see for example [15]) are formulated in terms of $\subseteq^{*}$ rather than $\subseteq$. The connection to the Tukey order is noticeable once the Tukey order is expressed in terms of cofinal maps. Given a partially ordered set $(D,\leq_{D})$, a set $B\subseteq D$ is called cofinal in $(D,\leq_{D})$ if for every $d\in D$ there is $b\in B$ such that $d\leq_{D}b$. Schmidt Duality [43] in the special case of ultrafilters (or if there are least upper bounds) asserts that: For any two ultrafilters $U,W$, $U\leq_{T}W$ if and only if there is a function $f:W\to U$ which is monotone (i.e. $A\subseteq B\Rightarrow f(A)\subseteq f(B)$) and has a cofinal image in $U$.
@@
It is not hard to check that whenever $U$ is a simple $P_{\mathbb{D}}$-point, then $(U,\supseteq^{*})\equiv_{T}\mathbb{D}$. So, the analysis in this paper is finer than the Tukey-type analysis of ultrafilters, but can be used to control it. For example, we use our construction to find an ultrafilter $U$ which is Tukey-equivalent to $\lambda_{0}\times\lambda_{1}$. From this we can analyze the Tukey spectrum and the depth spectrum (see Definitions 5.3, 5.9) and show the consistency of an ultrafilter with a non-convex spectrum (both depth and Tukey). This was asked by the first author in [10].
@@
$\mathfrak{b}_{\kappa}=\min\{|\mathcal{F}|\mid\mathcal{F}\subseteq\kappa^{\kappa}\text{ is }{\leq^{*}}\text{-unbounded}\}$
$\mathfrak{d}_{\kappa}=\min\{|\mathcal{F}|\mid\mathcal{F}\subseteq\kappa^{\kappa}\text{ is }{\leq^{*}}\text{-dominating}\}$
@@
As we mentioned before, at a measurable cardinal, the only previously known method to produce models where $\mathfrak{u}_{\kappa}<2^{\kappa}$ is via the linear Mathias iteration. The computation of the values of these cardinal characteristics was done in [16] where it was shown that after a linear iteration of Mathias forcing, $\mathfrak{u}_{\kappa}=\mathfrak{u}^{com}_{\kappa}=\mathfrak{b}_{\kappa}=\mathfrak{d}_{\kappa}$. Lately, the first and the third author [7] generalized this and showed that merely the presence of a simple $P_{\lambda}$-point (which is the kind of ultrafilters produced by the Mathias iteration) causes $\mathfrak{u}_{\kappa}=\mathfrak{u}^{com}_{\kappa}=\mathfrak{b}_{\kappa}=\mathfrak{d}_{\kappa}=\lambda$. So the question of separating the cardinal characteristics in a model where $\mathfrak{u}_{\kappa}<2^{\kappa}$ remained open. Using our methods, we can prove the following:
@@ 2. Some preliminaries
Given a partial order $\mathbb{D}=(D,\leq_{D})$, we define $d_{1}<_{D}d_{2}$ if $d_{1}\leq_{D}d_{2}$ and $d_{1}\neq d_{2}$. For each $d\in D$, let $D_{<d}=\{e\in D:e<_{D}d\}$; $D_{\leq d}$ is defined similarly.
@@
Suppose otherwise, that $\mathbb{D}=(D,\leq_{D})$ and let $D=\bigcup_{i<\kappa}D_{i}$, where ${\langle}D_{i}\mid i<\kappa{\rangle}$ is an $\subseteq$-increasing sequence of $\kappa$-directed non-cofinal subsets of $D$. By moving to a subsequence of the $D_{i}$’s if necessary, we may assume that for every $i$, there is an element $e_{i}$ of $D_{i+1}$ which is not bounded by any element of $D_{i}$.
@@
Since $X_{\text{even}}\cup X_{\text{odd}}=A^{*}\in U$, either $X_{\text{even}}\in U$ or $X_{\text{odd}}\in U$. Suppose for example that $X_{\text{even}}\in U$ (the argument in the case that $X_{\text{odd}}\in U$ is identical). Then there is $i_{1}<\kappa$ such that $X_{\text{even}}\in F_{i_{1}}$. By $(2)$ of the claim, and $\kappa$-completeness of $F_{i_{1}}$, $\bigcap_{\alpha<i_{1}}A_{\alpha}\in F_{i_{1}}$. Therefore $X_{\text{even}}\cap\bigcap_{\alpha<i_{1}}A_{\alpha}\in F_{i_{1}}$. Since the $F_{r}$’s are increasing, we may assume that $i_{1}=2\gamma+1$ for some $\gamma<\kappa$. Finally, we claim that $X_{\text{even}}\cap\bigcap_{\alpha<2\gamma+1}A_{\alpha}\subseteq A_{2\gamma+1}$. This produces a contradiction since $X_{\text{even}}\cap\bigcap_{\alpha<2\gamma+1}A_{\alpha}\in F_{2\gamma+1}$ while $A_{2\gamma+1}\notin F_{2\gamma+1}$. Indeed if $\nu\in X_{\text{even}}\cap\bigcap_{\alpha<2\gamma+1}A_{\alpha}$ then by definition of $X_{\text{even}}$ there is $\alpha^{*}<\kappa$ such that $\nu\in C_{2\alpha^{*}}$. Since $\nu\in\bigcap_{\alpha<2\gamma+1}A_{\alpha}$, it follows that $2\alpha^{*}>2\gamma+1$. This means that $\nu\in A_{2\gamma+1}.$
@@
The ultrafilters we will be constructing are all $P$-points. This poses more restrictions on the possible $\mathbb{D}$ which we will be able to realize.
Recall that an ultrafilter $U$ is a $P_{\lambda}$-point, if $(U,\supseteq^{*})$ is $\lambda$-directed.
@@
3. $\mathbb{D}$-based iteration
In this section we will describe a forcing iteration which is indexed by a well-founded partial order. As we noted in the introduction, similar iterations appeared in various places in the literature. Notably, in the works of Shelah (see, for example, [44, 45]), this type of iteration is modeled as linear iteration with restricted memory — so in Shelah’s language, asking the memory to be transitive is parallel to forcing along a partial order. In [45], Shelah described a more general mechanism that allows one to deal with certain schemes which might be ill–founded or non-transitive.
We describe the general method with respect to $\kappa-$supports where $\kappa$ is either $\aleph_{0}$ (so, in this case we are referring to finite support) or measurable. As the treatment of the basic properties of the forcing in both cases is very similar, we include both of them in this section.
Let us define the notion of $\kappa-$support iteration based on $\mathbb{D}$ uniformly for every partial order $\mathbb{D}$.
Let $\mathbb{D}=(D,\leq_{D})$ be a partial order and let $\kappa$ be a regular cardinal, a sequence $\langle\mathbb{P}_{<d},\dot{\mathbb{Q}}_{d}:d\in D\rangle$ is a $\kappa-$support iteration based on $\mathbb{D}$ if the following conditions hold for all $d\in D$:
  • $\mathbb{P}_{<d}$ is the set of $\kappa-$supported sequences $p=\langle p_{e}\mid e\in D_{<d}\rangle$ such that for all $e\in D_{<d}$, $\Vdash_{\mathbb{P}_{<e}}p(e)\in\dot{\mathbb{Q}}_{e}$.
  • $\Vdash_{\mathbb{P}_{<d}}\dot{\mathbb{Q}}_{d}\text{ is a partial order}$.
We denote by $\mathbb{P}_{\mathbb{D}}$ the set of $\kappa-$supported sequences $p=\langle p_{e}\mid e\in D\rangle$ such that for all $e\in D$, $\Vdash_{\mathbb{P}_{<e}}p(e)\in\dot{\mathbb{Q}}_{e}$. The order is defined recursively by
$p\leq_{\mathbb{P}_{\mathbb{D}}}q\text{ iff for every }d\in D,\ p\restriction D_{<d}\Vdash_{\mathbb{P}_{<d}}p(d)\leq_{\dot{\mathbb{Q}_{d}}}q(d).$
We will also write $\mathbb{P}_{\leq d}$ instead of $\mathbb{P}_{<d}*\dot{\mathbb{Q}}_{d}$. Ultimately, we will only consider iterations based on well-founded partially ordered sets, as the names $\dot{\mathbb{Q}}_{d}$ will be defined by well-founded recursion.
Lemma 3.1.
Let $\mathbb{D}^{\prime}\subseteq\mathbb{D}$ be downwards closed. Then $\mathbb{P}_{\mathbb{D}}$ projects via restriction to $\mathbb{P}_{\mathbb{D}^{\prime}}.$